Consider a sample of calcium carbonate in the form of a cube measuring 2.205 in. on each edge.
Part A
If the sample has a density of 2.72 g/cm3 , how many oxygen atoms does it contain?
1
Expert's answer
2019-09-30T05:14:53-0400

Density of the sample"=2.72\frac{g}{cm^3}"

Edge length"=2.205in=2.205\times2.54=5.6007cm"

So the volume of the cubic sample"=5.6007\times5.6007\times5.6007=175.68cm^3"

Hence the mass of the sample"=2.72\times175.68=477.85g"

Molar mass of "CaCO_3" "=100\frac{g}{mol}"

Therefore number of moles of "CaCO_3" in the given sample"\frac{477.85}{100}=4.7785mol"

In "1mol" "CaCO_3", number of "O" atoms"=3\times N_A"

So in "4.7785mol" "CaCO_3", number of "O" atoms"=4.7785\times3\times N_A=8.634\times10^{24}atoms"